# Unified Taxonomy for Multivariate Time Series Anomaly Detection using Deep Learning

### 19 Mar 2026 v1

@@ Abstract
The topic of Multivariate Time Series Anomaly Detection (MTSAD) has grown rapidly over the past years, with a steady rise in publications and Deep Learning (DL) models becoming the dominant paradigm. To address the lack of systematization in the field, this study introduces a novel and unified taxonomy with eleven dimensions over three parts (Input, Output and Model) for the categorization of DL-based MTSAD methods. The dimensions were established in a two-fold approach. First, they derived from a comprehensive analysis of methodological studies. Second, insights from review papers were incorporated. Furthermore, the proposed taxonomy was validated using an additional set of recent publications, providing a clear overview of methodological trends in MTSAD. Results reveal a convergence toward Transformer-based and reconstruction and prediction models, setting the foundation for emerging adaptive and generative trends. Building on and complementing existing surveys, this unified taxonomy is designed to accommodate future developments, allowing for new categories or dimensions to be added as the field progresses. This work thus consolidates fragmented knowledge in the field and provides a reference point for future research in MTSAD.
@@ 1 Introduction
Multivariate Time Series Anomaly Detection (MTSAD) has gained relevant attention in recent years, reflected by a growing interest of the scientific community and its widespread adoption across domains such as industrial systems [WADI], server monitoring [OmniAnomaly] and spacecraft operations [MSL]. MTSAD approaches offer clear advantages over Univariate Time Series (UTS) methods [BlzquezGarca2021], as MTSAD involves capturing both the temporal dependencies typical of time series data and the intermetric (aka spatial) dependencies that reflect relationships among variables. This spatio-temporal nature of MTSAD approaches enables the detection of distinct anomaly patterns across one, several or all variables simultaneously, which enhances the identification of complex and subtle anomalies [Fu2024, Guan2024, He2020, Tian2023, MSCRED, MTAD-GAT]. Given the complexity of MTSAD task, conventional statistical methods like Vector Autoregressive (VAR) models and classical Machine Learning (ML) approaches, such as K-means clustering and Isolation Forest (IF) [Audibert2022] originally designed for UTS, fall short in addressing its challenges. In contrast, Deep Learning (DL) models have emerged as state-of-the-art solutions, due to their ability to model non-linear temporal dynamics and dependencies across multiple variables that traditional methods often miss.
@@
An important dimension to consider is how DL models capture temporal and spatial dependencies. For instance, Choi2021 emphasize this distinction and ZamanzadehDarban2024 categorize methods as Temporal (T), Spatial (S) or Spatio-temporal (ST). Li2023 further discuss thresholding techniques, feature extraction and evaluation metrics. Despite these valuable contributions, the DL-based MTSAD field still lacks a unified framework. While several reviews cover similar dimensions, they often propose independent and sometimes inconsistent categorizations. Moreover, key aspects remain underexplored, e.g. how spatial and temporal dependencies are jointly modeled in spatio-temporal approaches. To address these gaps and bring more clarity to the field, this paper proposes a taxonomy of DL-based models aiming MTSAD. The taxonomy has eleven key dimensions grouped into three main parts, to characterize Input, Output and Model. Building on concepts from recent reviews [Choi2021, Correia2024, Li2023, ZamanzadehDarban2024] and enriched by insights from an extensive screening of literature published between 2019 and 2024 and indexed at Scopus®, (https://www.scopus.com/), the taxonomy organizes MTSAD methods into a coherent and unified framework. The new taxonomy not only facilitates systematic comparison across existing methods but also highlights opportunities for future research. It aims to serve as a practical guide for researchers and practitioners, as well as a starting point for newcomers to the field. In the near future, the taxonomy is expected to be expanded and refined alongside advances in model architectures, detection strategies and data representations.
The remainder of this paper is organized as follows. Section 2 introduces the key concepts, problem formulation, performance evaluation methodologies for MTSAD models and benchmarks. Section 3 describes the screening process used to select relevant papers for this review and presents the proposed taxonomy of DL-based approaches for MTSAD. Section 4 provides a visual summary of the literature distribution across the key dimensions and discusses representative approaches within each dimension of the taxonomy, highlighting their design choices, strengths, and limitations. Section 5 examines the current research paradigm, identifies prevailing patterns, and provides an overview of recent developments, ongoing challenges, and potential future research directions. Finally, Section 6 concludes the paper.
@@ 2.1 Types of anomalies: temporal versus inter-metric
When dealing with MTS data, anomalies are likely to arise along the temporal dimension (temporal anomaly) and across different metrics at a given time point (inter-metric anomaly). While a temporal anomaly is typically defined along the temporal axis for each metric (e.g. Point, Contextual or Collective anomaly), the inter-metric anomaly occurs when the relationship between two or more metrics changes within a specific time interval, even if the individual metrics appear normal when considered separated. In practice, a temporal anomaly in a subset of metrics often corresponds to a temporal inter-metric anomaly, since deviations in one metric may affect the correlation with other metrics. Therefore, anomalies in MTS can manifest both temporally and inter-metrically, which require detection methods that capture the temporal dynamics along each metric as well as the relationship across metrics. Figure 1 shows representative examples of anomalies in MTS data. A temporal anomaly typically manifests when the evolution of a single metric deviates from its historical or expected temporal pattern, such as an abrupt spike (Point anomaly, A), an unexpected deviation in a given context (Contextual anomaly, B) or a subsequence of values that collectively exhibit abnormal behavior (Collective anomaly, C and D). In contrast, an inter-metric anomaly occurs when the statistical or functional relationship between two or more metrics changes, even if the individual metrics appear normal in isolation. This is the e.g. case for the anomaly (C) in Fig. 1 which is also inter-metric as it disrupts the dependency structure between metrics.
@@ 2.2 Problem Formulation in MTSAD
that represents the information at time point $t=1,\cdots,T$ over all $m=1,\cdots,M$ metrics. Alternatively, $\mathbf{X}:=\textbf{x}_{\bullet,t:t+w}\in\mathbb{R}^{M\times(w+1)}$ can be a subsequence (or window) of length $w+1$ denoted as
@@
which captures both short-term temporal patterns and associations across metrics. This unified notation allows the framework to detect Contextual and Collective anomalies besides Point ones, via window-based detection. In certain settings, the raw time series may be further transformed into alternative representations, such as descriptive measures based on sufficient statistics, learned embeddings or even images, that can also serve as input to the model.
In classification settings, the model directly provides a binary output $\hat{y}=f(\mathbf{X})$ given the input $\mathbf{X}$. On the other hand, a regression model produces an output $\hat{\mathbf{X}}=f(\mathbf{X})$, which can be a representation (embedding), reconstruction or prediction based on the input data. The anomaly score is then computed as the dissimilarity between the observed $\mathbf{X}$ and the model output $\hat{\mathbf{X}}$,
@@ 2.4 Benchmark MTSAD methodologies and datasets
Comparing MTSAD methodologies is challenging due to variations in datasets, evaluation metrics and experimental setups. Several real-world data benchmarks have been widely adopted, including spacecraft datasets (SMAP and MSL [MSL]), water treatment and distribution systems (SWaT [SWAT] and WADI [WADI]) and server monitoring environments (SMD [OmniAnomaly] and PSM [Rancoders]). Synthetic datasets, such as NeurIPS-TS [Lai2021], offer controlled anomaly rates and types, facilitating systematic evaluation of model behavior under different conditions. See other examples of data benchmarks in Table 3 of [ZamanzadehDarban2024].
@@ 3 Taxonomy of DL-based approaches for MTSAD
This paper proposes a taxonomy for DL in MTSAD to provide a structured organization of the state-of-the-art, unifying the results from the analysis of 409 documents and 12 reviews or evaluation papers. The following sections present the paper screening process and selection criteria, publication trends in the research field and the proposed taxonomy in detail. The full list of selected documents and reviews is available at the supplementary material of this paper.
@@ 3.3 The proposed taxonomy on MTSAD
Input Pre-processing Yes | No Raw TS | Transformed TS
@@
Optimal Thresholding Yes | No | Not Applicable Data-driven thresholding strategy
Time-Varying Threshold Yes | No | Not Applicable Threshold adapts over time
Granularity Point / Subsequence Resolution of anomaly labeling
@@ 3.4 Insights from prior research
The dimensions of the taxonomy were derived from the analysis of the screened documents and complemented by insights from review papers [Choi2021, Correia2024, Li2023, ZamanzadehDarban2024]. This approach allowed to derive a unified taxonomy upon prior work while refining it to reflect the current state-of-the-art. Regarding the Input part, ZamanzadehDarban2024 proposed a categorization focused on the Format dimension, distinguishing between Point and Window; however, screening methodological papers revealed these categories are insufficient to fully characterize MTSAD inputs, motivating the addition of an Image category to Format and the introduction of the new dimensions Preprocessing and Content. The dimensions of the Output part were also defined based on insights from prior reviews [Choi2021, Li2023, ZamanzadehDarban2024]. Choi2021 groups the methodologies based on how the anomaly score is computed (either via reconstruction error, prediction error or dissimilarity) while ZamanzadehDarban2024 categorized as forecasting-based, reconstruction-based, representation-based or hybrid. The dissimilarity category broadly aligns with the representation-based category, as both rely on comparisons in transformed feature spaces. Therefore, the categories from ZamanzadehDarban2024 were adopted for the Type dimension. For the Threshold dimension, Li2023 proposed an extensive taxonomy including fixed, optimal, adaptive, specific and non-parametric thresholds. However, this scheme mixes characteristics with computation methods and includes overlapping categories. The current taxonomy simplifies it by proposing two new binary dimensions, Optimal Threshold and Time-varying Threshold. Lastly, the Granularity dimension captures the level at which the anomaly detection output is produced, distinguishing between individual time points and (possibly overlapping) sequences of time steps. This differs from the categorization of anomaly types, where (point, subsequence) [ZamanzadehDarban2024] or (point, interval, series) [Li2023] describe the anomaly type rather than the output resolution.
@@ 4 Overview of the Literature on the Dimensions of the New Taxonomy
Figure 6 presents the classification of the 409 categorized models according to the proposed taxonomy. The three Parts of the taxonomy are shown at the innermost circle of the Sunburst. Moving outward, the eleven dimensions of the taxonomy are displayed and the outermost layer corresponds to the specific categories within each dimension. Each model is assigned to exactly one category per dimension. Colors in the chart indicate the number of documents assigned to each category, providing a visual sense of their relative prevalence. For some dimensions, such as Loss Function, Time-varying Threshold and Optimal Threshold, certain papers did not provide explicit information; these cases are represented by the category labeled “n.e.” (not explicit). The following sections provide a more detailed overview of the literature aligned with each dimension of the new taxonomy.
Figure 6: Sunburst chart of 409 categorized documents across the taxonomy. The inner circle shows the Parts, followed by Dimensions and Categories. Colors represent the number of documents and “n.e.” (not explicit) refers to papers that did not explicitly report information for a given dimension. The iterative version of this figure is available at https://github.com/ieeta-pt/DL-MTSAD.
@@ 4.1 Input: Preprocessing, Format and Content
The dimensions of the Input include Preprocessing, Format and Content. Regarding Content (Time | Others | Time + Others), it is clear that most studies preserve and make use of the temporal ordering of the time series, reflecting its fundamental importance for anomaly detection. Temporal data is typically used independently, and just a subset of studies enhances it with additional attributes such as spatial correlations, temporal dependencies or frequency-driven features. Some studies focus exclusively on non-temporal information, with spatial correlations being the most common, highlighting the role of relationships among variables in multivariate anomaly detection. In terms of Format (Point | Window | Image | Window + Image), the majority of inputs are structured as Window, which help preserve temporal context and facilitate the detection of Collective and Contextual anomalies, whereas only a few studies make use of time Point. Although most studies utilize a sliding window technique to extract fixed-length windows, temporal segmentation can alternatively be applied, resulting in non-overlapping segments with different lengths (e.g. seq2GMM). Moreover, some studies convert MTS windows into Image representations to capture temporal or spatial correlations, and just a few combine both Window and Image formats to enrich the input with complementary information. With respect to Preprocessing (Yes | No), most studies do not apply any preprocessing step and instead rely directly on raw time series as input. When preprocessing is applied, it typically emphasizes temporal characteristics of the data through filtering, decomposition or multi-scale transformations [SDDformer, Meta-MWDG, TFAD]. Other strategies focus on frequency-based representations or on explicitly computing spatial and temporal correlations, either alone or combined with raw series, to enrich the input of the DL-model [STFT-TCAN].
@@
$r_{i,j}=\frac{\sum_{\delta=0}^{w}x_{i}^{t-\delta}x_{j}^{t-\delta}}{\kappa},\quad i,j=1,2,...,M\ ,$ (13)

### 24 Apr 2026 v2

@@ Abstract
The topic of Multivariate Time Series Anomaly Detection (MTSAD) has grown rapidly over the past years, with a steady rise in publications and Deep Learning (DL) models becoming the dominant paradigm. To address the lack of systematization in the field, this study introduces a novel and unified taxonomy with eleven dimensions over three parts (Input, Output and Model) for the categorization of DL-based MTSAD methods. The dimensions were established in a two-fold approach. First, they derived from a comprehensive analysis of methodological studies. Second, insights from review papers were incorporated. Furthermore, the proposed taxonomy was validated using an additional set of recent publications, providing a clear overview of methodological trends in MTSAD. Results reveal a convergence toward Transformer-based and reconstruction models, setting the foundation for emerging adaptive and generative trends. Building on and complementing existing surveys, this unified taxonomy is designed to accommodate future developments, allowing for new categories or dimensions to be added as the field progresses. This work thus consolidates fragmented knowledge in the field and provides a reference point for future research in MTSAD.
@@ 1 Introduction
Multivariate Time Series Anomaly Detection (MTSAD) has gained relevant attention in recent years, reflected by a growing interest of the scientific community and its widespread adoption across domains such as industrial systems [WADI], server monitoring [OmniAnomaly] and spacecraft operations [MSL]. Unlike Univariate Time Series (UTS) anomaly detection methods [BlzquezGarca2021], which model only temporal dependencies within a single series, MTSAD requires capturing both temporal and inter-metric (i.e., spatial) dependencies that reflect relationships among variables. This results in a more complex modeling task, as methods must account for both intra- and inter-series dynamics. When such multivariate dependencies are present, this spatio-temporal setting allows the identification of a broader range of anomaly patterns that may affect one, several, or all variables simultaneously [Fu2024, Guan2024, He2020, Tian2023, MSCRED, MTAD-GAT]. Given the complexity of MTSAD task, conventional statistical methods like Vector Autoregressive (VAR) models and classical Machine Learning (ML) approaches, such as K-means clustering and Isolation Forest (IF) [Audibert2022] originally designed for UTS, fall short in addressing its challenges. In contrast, Deep Learning (DL) models have emerged as state-of-the-art solutions, due to their ability to model non-linear temporal dynamics and dependencies across multiple variables that traditional methods often miss.
@@
An important dimension to consider is how DL models capture temporal and spatial dependencies. For instance, Choi2021 emphasize this distinction and ZamanzadehDarban2024 categorize methods as Temporal (T), Spatial (S) or Spatio-temporal (ST). Li2023 further discuss thresholding techniques, feature extraction and evaluation metrics. Despite these valuable contributions, the DL-based MTSAD field still lacks a unified framework. While several reviews cover similar dimensions, they often propose independent and sometimes inconsistent categorizations. Moreover, key aspects remain underexplored, e.g. how spatial and temporal dependencies are jointly modeled in spatio-temporal approaches.
To address these gaps, this paper proposes a unified taxonomy that, unlike existing surveys, systematically covers the key aspects of DL-based models targeting MTSAD. Built on concepts from recent reviews [Choi2021, Correia2024, Li2023, ZamanzadehDarban2024] and enriched by insights from an extensive screening of literature published between 2019 and 2024 indexed at Scopus® (https://www.scopus.com/), it comprises eleven key dimensions grouped into three main parts characterizing Input, Output and Model. This provides a consistent and extensive framework that enables direct comparison across methods and explicitly addresses aspects that have so far remained underexplored, highlighting opportunities for future research. This novel taxonomy aims to serve as a practical guide for researchers and practitioners, as well as a starting point for newcomers to the field. In the near future, the taxonomy is expected to be expanded and refined alongside advances in model architectures, detection strategies and data representations.
This paper is organized in a crescendo of detail and complexity. Section 2 serves as a background for readers entering the research field, providing key concepts, problem formulation, performance evaluation methodologies, and benchmarks for MTSAD. Section 3 describes the methodology used to construct the taxonomy, starting with the paper screening and selection process and culminating in the proposed taxonomy of DL-based approaches for MTSAD, complemented by insights gathered from previous reviews. Section 4 presents a visual summary of the literature distribution across the taxonomy dimensions and discusses representative approaches within each dimension, highlighting their design choices, strengths, and limitations. Section 5 examines the current research paradigm, identifies prevailing patterns, and outlines recent developments, ongoing challenges, and future research directions. Finally, Section 6 concludes the paper.
@@ 2.1 Types of anomalies: temporal versus inter-metric
When dealing with MTS data, anomalies are likely to arise along the temporal dimension (temporal anomaly) and across different metrics at a given time point (inter-metric anomaly). While a temporal anomaly is typically defined along the temporal axis for each metric (e.g. Point, Contextual or Collective anomaly) [Chandola2009], the inter-metric anomaly occurs when the relationship between two or more metrics changes within a specific time interval, even if the individual metrics appear normal when considered separated [Interfusion]. In practice, a temporal anomaly in a subset of metrics often corresponds to a temporal inter-metric anomaly, since deviations in one metric may affect the correlation with other metrics. Therefore, anomalies in MTS can manifest both temporally and inter-metrically, which require detection methods that capture the temporal dynamics along each metric as well as the relationship across metrics. Figure 1 shows representative examples of anomalies in MTS data. A temporal anomaly typically manifests when the evolution of a single metric deviates from its historical or expected temporal pattern, such as an abrupt spike (Point anomaly, A), an unexpected deviation in a given context (Contextual anomaly, B) or a subsequence of values that collectively exhibit abnormal behavior (Collective anomaly, C and D) [Chandola2009]. In contrast, an inter-metric anomaly occurs when the statistical or functional relationship between two or more metrics changes, even if the individual metrics appear normal in isolation [Interfusion]. This is the e.g. case for the anomaly (C) in Fig. 1 which is also inter-metric as it disrupts the dependency structure between metrics.
@@ 2.2 Problem Formulation in MTSAD
that represents the information at time point $t=1,\cdots,T$ over all $m=1,\cdots,M$ metrics [Rancoders]. Alternatively, $\mathbf{X}:=\textbf{x}_{\bullet,t:t+w}\in\mathbb{R}^{M\times(w+1)}$ can be a subsequence (or window) of length $w+1$ denoted as
@@
which captures both short-term temporal patterns and associations across metrics [OmniAnomaly, USAD]. This unified notation allows the framework to detect Contextual and Collective anomalies besides Point ones, via window-based detection. In certain settings, the raw time series may be further transformed into alternative representations, such as descriptive measures based on sufficient statistics, learned embeddings or even images, that can also serve as input to the model [MSCRED].
In classification settings, the model directly provides a binary output $\hat{y}=f(\mathbf{X})$ given the input $\mathbf{X}$ [MBead]. On the other hand, a regression model produces an output $\hat{\mathbf{X}}=f(\mathbf{X})$, which can be a representation (embedding) [DCdetector], reconstruction [OmniAnomaly] or prediction [GDN] based on the input data. The anomaly score is then computed as the dissimilarity between the observed $\mathbf{X}$ and the model output $\hat{\mathbf{X}}$,
@@ 2.4 Benchmark MTSAD methodologies and datasets
MTSAD methods have applications across a diversity of real-world domains, including industrial systems such as power plants [MSCRED, Choi2020], IoT and cyber-physical systems such as water distribution networks [GDN, MAD-GAN], robotics [SWCVAE, LSTM-VAE], server machine monitoring [OmniAnomaly, Rancoders], healthcare [CAE-M], space telemetry [MSL] and structural health monitoring [STVAE].
Comparing MTSAD methodologies is challenging due to variations in datasets, evaluation metrics and experimental setups. Several real-world data benchmarks have been widely adopted, including SMAP and MSL [MSL], SWaT [SWAT] and WADI [WADI] and SMD [OmniAnomaly] and PSM [Rancoders]. Synthetic datasets, such as NeurIPS-TS [Lai2021], offer controlled anomaly rates and types, facilitating systematic evaluation of model behavior under different conditions. See other examples of data benchmarks and their domains in Table 3 of [ZamanzadehDarban2024].
@@ 3 Taxonomy of DL-based approaches for MTSAD
This paper proposes a taxonomy for DL in MTSAD to provide a structured organization of the state-of-the-art, unifying the results from the analysis of 409 documents and 12 reviews or evaluation papers. The full list of categorized papers and reviews is available in the supplementary material of this paper. Figure 4 illustrates the paper screening process, resulting in 409 categorized papers from an initial pool of 898 screened documents. The following sections present the paper screening process and selection criteria, publication trends in the research field and the proposed taxonomy in detail.
Figure 4: Paper screening and selection process. Starting from 898 screened documents (2019–2024), 393 DL-based papers were selected and combined with 16 papers from reference list screening, resulting in 409 categorized papers. Additionally, 12 review papers were identified to complement the proposed taxonomy.
@@ 3.3 The proposed taxonomy on MTSAD
Figure 7: Schematic diagram of the Model dimensions (in bold). The dashed border indicates that the Feature Extraction module is optional.
@@
Input Preprocessing Yes | No Raw TS | Transformed TS
@@
Optimal
Threshold
Yes | No | Not Applicable Data-driven thresholding strategy
Time-varying
Threshold
Yes | No | Not Applicable Threshold adapts over time
Granularity Point | Subsequence Resolution of anomaly labeling
@@ 3.4 Insights from prior research
The dimensions of the taxonomy were derived from the analysis of the screened documents and complemented by insights from review papers [Choi2021, Correia2024, Li2023, ZamanzadehDarban2024]. This approach allowed to derive a unified taxonomy upon prior work while refining it to reflect the current state-of-the-art. Table 2 presents literature contributions mapped into the dimensions of the proposed taxonomy. Regarding the Input part, ZamanzadehDarban2024 proposed a categorization focused on the Format dimension, distinguishing between Point and Window; however, screening methodological papers revealed these categories are insufficient to fully characterize MTSAD inputs, motivating the addition of an Image category to Format and the introduction of the new dimensions Preprocessing and Content. The dimensions of the Output part were also defined based on insights from prior reviews [Choi2021, Li2023, ZamanzadehDarban2024]. Choi2021 groups the methodologies based on how the anomaly score is computed (either via reconstruction error, prediction error or dissimilarity) while ZamanzadehDarban2024 categorized as forecasting-based, reconstruction-based, representation-based or hybrid. The dissimilarity category broadly aligns with the representation-based category, as both rely on comparisons in transformed feature spaces. Therefore, the categories from ZamanzadehDarban2024 were adopted for the Type dimension. For the Threshold dimension, Li2023 proposed an extensive taxonomy including fixed, optimal, adaptive, specific and non-parametric thresholds. However, this scheme mixes characteristics with computation methods and includes overlapping categories. The current taxonomy simplifies it by proposing two new binary dimensions, Optimal Threshold and Time-varying Threshold. Lastly, the Granularity dimension captures the level at which the anomaly detection output is produced, distinguishing between individual time points and (possibly overlapping) sequences of time steps. This differs from the categorization of anomaly types, where (point, subsequence) [ZamanzadehDarban2024] or (point, interval, series) [Li2023] describe the anomaly type rather than the output resolution.
@@
Table 2: Literature contributions to the novel of taxonomy proposed in this work. — indicates it is absent or not explicitly defined.
Part Dimension Choi2021 Li2023 ZamanzadehDarban2024 Correia2024
Input Preprocessing
Format Point, Window Sliding window, Input sequence
Content
Output Type Reconstruction, Prediction, Dissimilarity Forecasting, Reconstruction, Representation, Hybrid Prediction, Reconstruction
Optimal
Threshold
Fixed, Optimal, Adaptive, Specific, Non-parametric
Time-variyng
Threshold
Granularity Point, Interval, Series Point, Subsequence
Model Task
Loss
Function
Adversarial, Reconstruction, Prediction, NLL
T/S
Dependency
Temporal, Spatial, Spatio-temporal
Feature
Extraction
Binary (yes/no)
@@ 4 Overview of the Literature on the Dimensions of the New Taxonomy
Figure 8 presents the classification of the 409 categorized models according to the proposed taxonomy. The three Parts of the taxonomy are shown at the innermost circle of the Sunburst. Moving outward, the eleven dimensions of the taxonomy are displayed and the outermost layer corresponds to the specific categories within each dimension. Each model is assigned to exactly one category per dimension. Colors in the chart indicate the number of documents assigned to each category, providing a visual sense of their relative prevalence. For some dimensions, such as Loss Function, Time-varying Threshold and Optimal Threshold, certain papers did not provide explicit information; these cases are represented by the category labeled “n.e.” (not explicit). An interactive version of this figure is available at https://ieeta-pt.github.io/DL-MTSAD/Sunburst.html, allowing readers to consult the counts per category and to explore the specific parts and dimensions of the taxonomy. The following sections provide a more detailed overview of the literature aligned with each dimension of the taxonomy, with representative examples of each category cited throughout. The complete list of categorized papers is also available at https://github.com/ieeta-pt/DL-MTSAD.
Figure 8: Sunburst chart of 409 categorized documents across the taxonomy. The inner circle shows the Parts, followed by Dimensions and Categories. Colors represent the number of documents and “n.e.” (not explicit) refers to papers that did not explicitly report information for a given dimension. The interactive version of this figure is available at https://ieeta-pt.github.io/DL-MTSAD/Sunburst.html.
@@ 4.1 Input: Preprocessing, Format and Content
The dimensions of the Input include Preprocessing, Format and Content. Regarding Content (Time | Others | Time + Others), it is clear that most studies preserve and make use of the temporal ordering of the time series, reflecting its fundamental importance for anomaly detection. Temporal data is typically used independently, and just a subset of studies enhances it with additional attributes such as spatial correlations, temporal dependencies or frequency-driven features. Some studies focus exclusively on non-temporal information, with spatial correlations being the most common, highlighting the role of relationships among variables in multivariate anomaly detection. Temporal-only inputs are simpler and computationally efficient, but may fail to capture inter-metric anomalies, whereas incorporating spatial or hybrid content can improve detection performance at the cost of increased model complexity and potential noise, when the additional content is weakly informative. In terms of Format (Point | Window | Image | Window + Image), the majority of inputs are structured as Window, which preserve temporal context and are necessary for the detection of Collective and Contextual anomalies, whereas only a few studies make use of time Point, which are suitable to detect Point anomalies. Although most studies utilize a sliding window technique to extract fixed-length windows, temporal segmentation can alternatively be applied, resulting in non-overlapping segments with different lengths (e.g. seq2GMM). Moreover, some studies convert MTS windows into Image representations to capture temporal or spatial correlations, and just a few combine both Window and Image formats to enrich the input with complementary information. Image representations enable the use of models typically used for image analysis, such as CNN, but at the cost of increased preprocessing complexity and potential loss of temporal resolution.
With respect to Preprocessing (Yes | No), most studies do not apply any preprocessing step and instead rely directly on raw time series as input. When preprocessing is applied, it typically emphasizes temporal characteristics of the data through filtering, decomposition or multi-scale transformations [SDDformer, Meta-MWDG, TFAD]. Other strategies focus on frequency-based representations or on explicitly computing spatial and temporal correlations, either alone or combined with raw series, to enrich the input of the DL-model [STFT-TCAN]. While preprocessing can improve detection by highlighting relevant signal characteristics, it also introduces additional design choices and may hide anomalies that exist in the raw time series.
@@
$r_{i,j}=\frac{\sum_{\delta=0}^{w}x_{i,t-\delta}x_{j,t-\delta}}{\kappa},\quad i,j=1,2,...,M\ ,$ (13)